Grade 11Physical Chemistry

in a 13 l vessel initially followingreaction occur c+s2=cs2 by 12 g c,64g s2,76g cs2 at 1027 deg c then total pressure is 200R ,158R,100R,79R

Spandan Chatterjee
8 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago

To tackle the scenario you've presented, we need to analyze the reaction and the conditions described. The reaction is as follows:

C + S2 ⇌ CS2

Initially, we have 12 grams of carbon (C) and 64 grams of sulfur (S2). The molar masses of these elements are approximately:

  • Carbon (C): 12 g/mol
  • Sulfur (S2): 32 g/mol (16 g/mol per sulfur atom)
  • Carbon disulfide (CS2): 76 g/mol

Step 1: Calculate Moles of Reactants

First, let's determine how many moles of carbon and sulfur we have:

  • Moles of Carbon (C) = 12 g / 12 g/mol = 1 mol
  • Moles of Sulfur (S2) = 64 g / 32 g/mol = 2 mol

So, we start with 1 mole of carbon and 2 moles of sulfur in the reaction vessel.

Step 2: Understand the Reaction Progress

When carbon reacts with sulfur, they form carbon disulfide (CS2). The stoichiometry of the reaction indicates that 1 mole of C reacts with 1 mole of S2 to produce 1 mole of CS2. Therefore:

  • From 1 mole of C and 1 mole of S2, we can produce 1 mole of CS2.

Since we have 1 mole of C and 2 moles of S2, carbon is the limiting reactant. Thus, 1 mole of CS2 will be produced.

Step 3: Calculate the Total Number of Moles at Equilibrium

After the reaction is complete, the changes in the number of moles will be as follows:

  • Moles of C left = 1 - 1 = 0
  • Moles of S2 left = 2 - 1 = 1
  • Moles of CS2 produced = 1

At equilibrium, the total moles in the vessel is:

  • Total moles = 0 (C) + 1 (S2) + 1 (CS2) = 2 moles

Step 4: Calculate the Total Pressure in the Vessel

Using the Ideal Gas Law (PV = nRT), we can find the pressure of the system. The total number of moles is 2, the volume of the vessel is 13 L, and R (the ideal gas constant) is 0.0821 L·atm/(K·mol). The temperature is 1027 °C, which we convert to Kelvin:

  • T(K) = 1027 + 273.15 = 1300.15 K

Now, substituting these values into the Ideal Gas Law to find pressure (P):

P = (nRT) / V

P = (2 moles * 0.0821 L·atm/(K·mol) * 1300.15 K) / 13 L

Calculating this gives:

P ≈ 200.2 atm

Step 5: Interpreting the Results

From the calculations, we find that the total pressure in the vessel is approximately 200 R. The values you provided (200 R, 158 R, 100 R, 79 R) seem to indicate different potential scenarios or conditions. However, based on the reaction and the ideal gas calculations, the pressure aligns closely with the 200 R value. If other pressures are observed, it may indicate the influence of temperature variations or other factors that weren't accounted for in this initial analysis.

In summary, the key steps in solving this problem involved determining the moles of reactants, understanding the stoichiometry of the reaction, and applying the Ideal Gas Law to find the total pressure in the system. This process illustrates not only the chemical principles at play but also the importance of precise measurements in chemical reactions.